# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 6.2 A Saks Type Integral
As we suppose that the space $Z$ is Hausdorff in its Boccuto-Candeloro net convergence structure, the integral just presented is well defined (taking into account that $\mathcal{G}_{f,A}$ is directed). Taking also the Fleishcer Net Riemann integral gives more flexibility in defining the integral for each $A\in\mathcal{H}$, instead only for the whole space (as in [15]).
As in the previous subsection, we are not going to study this integral in detail, but just point some interesting aspects of it in connection to the pervious ones.
In fact, if $f$ is already Fleischer Net Riemann integrable on $A\in\mathcal{H}$, taking $\mathcal{G}_{f,A}=\{A\}$ produces the following:
@@
If $f:\Omega\rightarrow X$ is Fleischer Net Riemann integrable on $A\in\mathcal{H}$, it is Saks integrable on $A$ and the two integrals agree.
@@
By the choice indicated above, i.e $\mathcal{G}_{f,A}=\{A\}$ (which is trivially upwards directed by inclusion), it follows that the net $\{\int_{B}fd\mu\}_{B\in\mathcal{G}_{f,A}}$ is constant and, by $(v)$ of Definition 3.3, its limit exists and equals the Fleischer Net Riemann integral $\int_{A}fd\mu$, which finishes the proof. ∎
Also, by ways of comparison, we point out that [15] defines a similar integral (which we call the Fleischer integral) [^8] Not to be confused with the Fleischer Net Riemann integral. based on the Kolmogorov S-integral with $\mathcal{H}$ being a ring and $\mathcal{G}_{f,A}$ being an ideal of sets of this ring and independent of $A$ (i.e, he takes it dependent only on $f$, and the only domain of integration being $\Omega$). In this case, we have:
@@
Suppose that $\mathcal{H}$ is a ring of subsets of $\Omega$. For each $f:\Omega\rightarrow X$ and $A\in\mathcal{H}$, take $\mathcal{G}_{f,A}$ to be an ideal of sets. Then, if for each $B\in\mathcal{G}_{f,A}$, $f$ is Kolmogorov Integrable (See 3.2) on $B$, the Saks integral exists and coincides with the Fleischer integral.
We also note that, by the work of Boccuto and Riecan [9], we may obtain that the Saks integral contains also the Choquet integral of Riesz space-valued functions ([8]), and the Sipos integral for the case of non-negative functions. In fact, let $\mathcal{H}$ be a $\sigma$-algebra and $C:\mathcal{H}\rightarrow Y$ be a capacity (for this notion, see [8]). Fix $f:\Omega\rightarrow X=\overline{\mathbb{R}}^{+}$ a function such that:
@@
and define $u_{f}(t)=u(t)=C(\Sigma^{f}_{t})$. Note that $u:\mathbb{R}\rightarrow Y$ is a monotone function. In this context, we shall need the next result (see Section 3 of [8]):
@@
where the two last integrals are in the sense of the Kolmogorov S-integral, and which produced exactly the Choquet integral of $f$ (fixed above) in Definition 3.3 of [8] with respect to $C$. Therefore, we get, by using also Theorem 3.6 of [8]:
@@
This shows that this idea of Riemann integration derived from Saks’ original ideas is also fruitful for the case of non-additive set functions.
Some convergence theorems (without proofs) for this integral in a more restricted setting can be found in the article [20].
We finish now the presentation of this integral, and indeed of all others, and pass to the final comments.
7 Classification Schemes and (Various) Open Problems
The various integrals presented above all possesses strengths and weaknesses, specially with respect to which classes of functions and measures are admitted as possible ingredients of the procedure. As Kolmogorov [13] argued, a general definition of an integration procedure is unlikely to contain all existing ones, for various reasons. Nonetheless, the development above provides a reason to expect that each integration procedure we can find falls within one of the four categories now presented:
@@ 7 Classification Schemes and (Various) Open Problems
    Weak Type Integration, after Pettis [136] famous model.
Given the variety and diversity of integration procedures exhibited in the present work, such a classification may be premature, but can generate a way of organizing all the massive existing literature about integration theories. In this respect, an open problem, which is more of a research program carried forward already by such works as [4], [67], [60] and others, is:
@@
An affirmative response to both problems seems unlikely, specially as some integration procedures are not even comparable, but even this is a useful response to the general question: is there a "most general" procedure in each class ? And, if so, what are its limits ?
@@
    Is it possible to obtain (uniform) convergence theorems for the integrals above without requiring some specific convergence, only working with general properties of the filters/net involved, maybe with some minor additional condition to guarantee good continuity behaviors of the convergence ?
This is basically a question of the continuity of a (linear and isotone, sometimes) functional on spaces with general convergences and with specific properties generated by integration. That seems to require additional study of general convergence structures and many-valued topologies as is [137], [138], [139], [140] and references and further works therein. This seems to be related to the connection between integration in general convergence structures and Daniell integration in this context - continuity properties here are the key to obtain possible connections.
With respect to measures and functions, it seems directly possible, and in fact straightforward, to modify all the Riemmanian constructions above, not so much the Lebesgue ones, to interval functions in the sense of Henstock [67]. This provides, in particular, a non-linear version of the integration procedures above. The domain of the measures produces another challenge:
@@
    Considering more general domains, such as orthoalgebras, difference posets, (D)-lattices and other structures presented in [141], which results above are preserved and, more interestingly, which convergence properties of measures and integrals are possible to prove ?
This also relates to the general integration theory of Kusraev and Malyugin [101], which may serve as a guideline, specially with respect to extension theorems of measures.
Nothing has been said with respect to weak type integrals of Pettis [136] type. It is not difficult to prove that the Fleischer Net Riemann and Saks Type Integrals (but not the $S^{*}$-integral!), are of the weak type as in Sipos [142] weak type order integral, which is defined in terms of an "order dual" of a Riesz Space. Related to duality in general convergence (partially ordered groups and vector spaces) as in Chapter 4 of [138], it arises the following question:
  • 1.
    Can a general weak-type Pettis integration be developed in partially ordered net convergence structures in such a way that this theory is compatible with Riemann and Lebesgue type-integration procedures in these general convergences ?
A comparison with and study of [143] and [144] should be made in this case.
Lastly, we pose the following problem:
  • 1.
    The spaces of integrable functions $L^{1}$, with respect to each integration procedures, can be characterized as partially ordered vector spaces. Which convergence and lattice structures turns them in a Riesz space with "good" (order separability, for example) properties ?
The references [145] and [41] can be a first step in this direction for general convergences.
@@ Acknowledgement
This study was funded in part by the Coordination for the Improvement of Higher Education Personnel - CAPES - Funding Code 001. The author would like to thank Professors Francesco Russo (ENSTA - Paris), Christophe Gallesco (UNICAMP), David Fremlin (Exeter) and Lech Drewnowski (Poznan) for interesting conversations and remarks regarding integration theory.

### v2

@@ Abstract
As we suppose that the space $Z$ is Hausdorff in its net convergence structure, the integral just presented is well defined (taking into account that $\mathcal{G}_{f,A}$ is directed).
Contrary to Corollary 2.2.1 and Theorem 2.3, the Saks integral will not be, in general, linear. Nevertheless, some general properties can be established in this general setting.
Denote by $L_{S}(\mu)$ the space of all functions $f:\Omega\rightarrow X$ Saks integrable on each $A\in\mathcal{H}$, with respect to the (fixed) set function $\mu:\Omega\rightarrow Y$ in the context of a (fixed) integration structure.
Theorem 6.1 (Properties of the Saks integral).
The Saks integral has the following properties compatible with a partially ordered integral structure in Riesz spaces $X,Y,Z$:
  1. (i)
    Let $f\in L_{S}(\mu)$ such that $f\geq 0$ and $\mu\geq 0$. Then,
    $\int^{Sa}_{A}fd\mu\geq 0,$
    for each $A\in\mathcal{H}$, i.e, the Saks integral is a positive map on the space of integrable functions.
  2. (ii)
    Let $\mu\geq 0$ and $f,g\in L_{NR}(\mu)$ such that:
    $g\geq f.$
    Then,
    $\int^{Sa}_{A}gd\mu\geq\int^{Sa}_{A}fd\mu,$
    for each $A\in\mathcal{H}$, i.e, the Saks integral is an isotone map on $L_{NR}(\mu)$ equipped with the partial ordering inherited from $X$.
  3. (iii)
    Let $\mu\geq 0$ and $f\in L_{Sa}(\mu)$ such that $|f|\in L_{NR}(\mu)$, then:
    $|\int^{Sa}_{A}fd\mu|\leq\int^{Sa}_{A}|f|d\mu,$
    for each $A\in\mathcal{H}$.
Its also easy to see that the Saks integral is more general than the net Riemann integral: if $f$ is net Riemann integrable on $A\in\mathcal{H}$, taking $\mathcal{G}_{f,A}=\{A\}$ produces the following:
@@
If $f:\Omega\rightarrow X$ is net Riemann integrable on $A\in\mathcal{H}$, it is Saks integrable on $A$ and the two integrals agree.
@@
By the choice indicated above, i.e $\mathcal{G}_{f,A}=\{A\}$ (which is trivially upwards directed by inclusion), it follows that the net $\{\int_{B}fd\mu\}_{B\in\mathcal{G}_{f,A}}$ is constant and, by Definition 1.13, its limit exists and equals the net Riemann integral $\int_{A}fd\mu$, which finishes the proof. ∎
Also, by ways of comparison, we point out that Fleischer (1990) defines a similar integral (which we call the Fleischer integral) based on the Kolmogorov S-integral with $\mathcal{H}$ being a ring and $\mathcal{G}_{f,A}$ being an ideal of sets of this ring and independent of $A$ (i.e, he takes it dependent only on $f$, and the only domain of integration being $\Omega$). In this case, we have:
@@
Suppose that $\mathcal{H}$ is a ring of subsets of $\Omega$. For each $f:\Omega\rightarrow X$ and $A\in\mathcal{H}$, take $\mathcal{G}_{f,A}$ to be an ideal of sets. Then, if for each $B\in\mathcal{G}_{f,A}$, $f$ is S-integrable (See 2.1) on $B$, the Saks integral exists and coincides with the Fleischer integral.
As Shcherbin (1991) integral is the same as Fleischer’s but in the context of Riesz spaces and order convergence, the result above also covers his integration procedure.
As indicated in the beginning of the section, non-additive integration procedures are of special interest in the context of Saks integration. In this regard, by using some elements of the work of Boccuto and Riečan (2009), we may obtain that the Saks integral contains also the Choquet integral of Riesz space-valued functions (Boccuto and Riečan (2008)), and the Sipos integral for the case of non-negative functions. In fact, let $\mathcal{H}$ be a $\sigma$-algebra and $C:\mathcal{H}\rightarrow Y$ be a capacity (for this notion, see Boccuto and Riečan (2008)). Fix $f:\Omega\rightarrow X=\overline{\mathbb{R}}^{+}$ a function such that:
@@
and define $u_{f}(t)=u(t)=C(\Sigma^{f}_{t})$. Note that $u:\mathbb{R}\rightarrow Y$ is a monotone function. In this context, we shall need the next result (see Section 3 of Boccuto and Riečan (2008)):
@@
where the two last integrals are in the sense of the Kolmogorov S-integral, and which produced exactly the Choquet integral of $f$ (fixed above) in Definition 3.3 of Boccuto and Riečan (2008) with respect to $C$. Therefore, we get, by using also Theorem 3.6 of Boccuto and Riečan (2008):
@@
Some convergence theorems for this integral in a more restricted setting can be found in the article Shcherbin (1991) and Gould (1965).
With these comments, we stop the development of this last integral, and consider some conclusions and open problems based on the integration theories so far defined.
7 Conclusion: Classification Schemes and (other) Open Problems.
The various integrals presented above all possesses strengths and weaknesses, specially with respect to which classes of functions and measures are admitted as possible ingredients of the procedure. As Kolmogorov in Kolmogoroff (1930) argued, a general definition of an integration procedure is unlikely to contain all existing ones, for various reasons. Nonetheless, the development above provides a reason to expect that each integration procedure we can find falls within one of the following four categories:
@@
    Weak Type Integration, after Pettis (1938) famous model.
Given the variety and diversity of integration procedures exhibited in the present work, such a classification may be premature, but can generate a way of organizing all the massive existing literature about integration theories. In this respect, an open problem, which is more of a research program carried forward already by such works as Boccuto and Sambucini (1997), Henstock (1991), Sion (1973) and others, is:
@@
An affirmative response to both problems seems unlikely, specially as some integration procedures are not even comparable in principle, but even this is a useful response to the general question: is there a "most general" procedure in each class ? And, if so, what are its limits ? Possibly, the main classes of integration procedures presented in this work are the most general in some sense.
@@
    Is it possible to obtain (uniform) convergence theorems for the integrals above without requiring some specific convergence and lattice norms, only working with general properties of the filters/net involved, maybe with some minor additional condition to guarantee good continuity behaviors of the convergence ?
This is basically a question of the continuity of a (linear and isotone, sometimes) functional on spaces with general convergences and with specific properties generated by integration. That seems to require additional study of general convergence structures and many-valued topologies as is Höhle (2001), Beattie and Butzmann (2002), Gähler (1977), Gähler (1978) and references and further works therein. This seems to be related to the connection between integration in general convergence structures and Daniell integration in this context - continuity properties here are the key to obtain possible connections.
With respect to measures and functions, the Riemmanian constructions above, not so much the Lebesgue ones, admit interval functions integration in the sense of Henstock (1991). This provides, in particular, a non-linear version of the integration procedures above. The domain of the measures produces another challenge:
@@
    Considering more general domains, such as orthoalgebras, difference posets, (D)-lattices and other structures presented in Dvurečenskij and Pulmannová (2000), which results above are preserved and, more interestingly, which convergence properties of measures and integrals are possible to prove ? That is, developing a non-commutative integration in the Riemmanian sense is an open problem in general.
@@
8 Appendix: Uniform Approximation of Measurable Functions with Values in Riesz Spaces.
Using the results of Section 5.1, some interesting integrability results can be achieved for classes of (Riesz space-valued) functions that are uniform order limits of sequences of simple and/or countably valued functions and measurable in some way. There appear to be only limited results in this area, with main references being Section 4 of Pavlakos (1983) and Proposition 2.1 of Potocky (1977). To not interrupt the main flux of the text, we prove new results in a wider setting in this appendix.
In fact, we propose a generalization of the results of these two references to specific types of locally solid topological Riesz spaces. In this case, let $X$ be Riesz space with a topology induced by a (possibly uncountable) collection of Riesz $F$-norms $\{p_{l}\}_{l\in L}$ in such a way that the sets:
$\bigcap_{i=1}^{n}\{y:p_{l_{i}}(x-y)\leq\varepsilon\},$
which depends on $x\in X$, $n\in\mathbb{N}$, $\varepsilon>0$ and $\{l_{i}\}_{i=1}^{n}$, forms a base of neighborhoods of a (Hausdorff) topology of $X$. Its easy to show that such sets are solid, in such a way that the induced topology is indeed a Riesz space locally solid topology (see (Zaanen, 1983, p. 296)).
Now, we define the concept of a strong order unit for the previously defined class of Riesz space (see Pavlakos (1983)):
Definition 8.1 (Strong Order Unit).
We say that a Riesz $F$-norm (on a Riesz space $X$) $p:X\rightarrow[0,\infty)$ has a strong order unit $u>0$ if $p(x)\leq k$ implies $|x|\leq ku$, for each $k\in\mathbb{R}_{+}$ and $x\in X$. We then say that a collection of Riesz $F$-norms $\{p_{l}\}_{l\in L}$ has a strong order unit if every element has it.
With the help of these concepts we can now state and prove the following result (compare with the two references above):
Proposition 8.1 (Approximation by Countably-Valued Functions).
Let $(X,\leq,\tau)$ be a locally solid Riesz space with topology $\tau$ induced by a collection $\{p_{l}\}_{l\in L}$ of Riesz $F$-norms. Suppose that $X$ is first countable and has a strong order unit with respect to its collection of Riesz $F$-norms. Then, a Borel-measurable (with respect to the class of open sets in the order topology of $X$) function $f:\Omega\rightarrow X$ is a order uniform limit of countably valued functions.
Proof.
As $X$ is first countable, it is separable and we can find a countable collection $C=\{c_{k}\}_{k\in\mathbb{N}}\subseteq X$ such that:
$f(\Omega)\subseteq\overline{C}=X.$
Then, denote by $C_{n,k}$ the subbasis element of the metrizable topology of $X$ consisting of the closed ball with radius $\frac{1}{n}$ of the element $c_{k}$ of $C$, with respect to the Riesz $F$-seminorm that induces its metrizable topology (by the Birkhoff-Kakutani theorem), which we denote by $p^{*}$. In this case, we make a disjoint sequence of elements by taking:
$B_{n,1}=C_{n,1},\ B_{n,k}=C_{n,k}-\bigcup_{i<k}C_{n,i},k\geq 2;$
in such a way that $\Omega=\bigcup_{k\in\mathbb{N}}B_{n,k}$ for every $n\in\mathbb{N}$.
Now, define the following sequence of countably valued functions defined on $\Omega$ with values in $X$:
$f_{n}(t)=\sum_{k=1}^{\infty}c_{k}\mathbbm{1}(t)_{C_{n,k}},$
for which is valid that:
$p^{*}(f(\omega)-f_{n}(\omega))\leq\frac{1}{n},$
for any $(\omega,n)\in\Omega\times\mathbb{N}$.
Hence, by the existence of a strong order unit $u$,
$|f(\omega)-f_{n}(\omega)|\leq\frac{1}{n}u,$
for any $(\omega,n)\in\Omega\times\mathbb{N}$. This shows that $f_{n}$ converges order uniformly to $f$ on $\Omega$. ∎
With the same proof, we may stablish the metrizable case:
Proposition 8.2 (Approximation by Countably-Valued Functions - II).
Let $(X,\leq,\tau)$ be a metrizable locally solid Riesz space with topology $\tau$. Suppose that $X$ is separable and has a strong order unit. Then, a Borel-measurable function $f:\Omega\rightarrow X$ is a order uniform limit of countably valued functions.
As a final note, it would be interesting to know the maximal (i.e, sufficient and necessary) conditions imposed into the Riesz space $X$ such that interesting classes of (measurable) functions are uniform limits, or even quasi-uniform limits, of nets of elementary/simple functions.
@@ Acknowledgement
This study was funded in part by the Coordination for the Improvement of Higher Education Personnel - CAPES - Funding Code 001. The author would like to thank Professors Francesco Russo (ENSTA - Paris), Christophe Gallesco (UNICAMP), David Fremlin (Exeter), Lech Drewnowski (Poznan) and Alberto Castejón (Vigo) for interesting conversations and remarks regarding integration theory.